Brief Exercise 19-4 Dilts Company has a unit selling price of $660, variable costs per unit of $440, and fixed costs of $265,000. Compute the break-even point in units using (a) the mathematical equation and (b) unit contribution margin. (Round answers to 0 decimal places, e.g. 5,275.) (a) Break-even point units (b) Break-even point units Click if you would like to Show Work for this question: Open Show Work
Answer #1

a.Let units sold at breakeven be x

Hence Sales-Variable cost-Fixed cost=0

660x-440x-265000=0

x(660-440)=265000

x=265000/220

=1205 units(Approx).

2.contribution margin=Sales price-Variable cost

=(660-440)=$220 per unit

Hence breakeven point=Fixed cost/contribution margin

=265000/220

=1205 units(Approx).